Aweemho:

A particular plant grows 3.5 inches per month. How many centimeters is the plant growing per month?

Shadow:

Think you just need to convert 3.5 inches to cm

Shadow:

\[\frac{ 3.5 inches}{ 1 } \times \frac{ 2.54 cm }{ 1 inches }\]

Shadow:

Multiply across then cancel